OpenStudy (anonymous):

show integral of 0 to infinity f(u) is 1/s laplace transform

OpenStudy (ybarrap):

Do you know that \(\frak{L}\tt^{-1}[ F(s)G(s)]=f(t)*g(t)\)? Where "*" is convolution?

OpenStudy (anonymous):

no...

OpenStudy (ybarrap):

To show the Laplace of \( \tt \int^\infty_0 f(u)~ du = \dfrac{1}{s} \), let \(\tt F(s)=\dfrac{1}{s}\) and \(\tt G(s)=1\) in the formula above. Do you see that?

OpenStudy (anonymous):

yes, wheres the 1 c from g(s) coming from

OpenStudy (ybarrap):

What we are doing is showing that the \(inverse \) of F(s)G(s) is the integral and this is the same as saying that the Laplace of the integral is 1/s.

OpenStudy (anonymous):

inverse os 1/s=s inverse of 1....that = s

OpenStudy (ybarrap):

Let's take a different approach.

OpenStudy (ybarrap):

This is convolution

OpenStudy (ybarrap):

$$f(t)*g(t) = \int^t_0 g(t-u)f(u) du$$ Have you seen this before?

OpenStudy (anonymous):

looks a little like product rule

OpenStudy (ybarrap):

Very different than the chain rule for derivatives, if that's what you're talking about. Convolution is what happens when you have a linear time-invariant system and you apply an input, here the input is g(t) and the system is f(t). So the operation of "integration" can be a 'system'. So the convolution says what happens in time (t) as you apply an input. For our system, it integrates the input. That's all convolution is. For different t, you will get a different value in the convolution, but all that's happening is summing (i.e. integrating).

OpenStudy (anonymous):

i have 1/s*1=integral 0 to infinity of 1 * t dt

OpenStudy (anonymous):

i think I'm understanding the theory

OpenStudy (anonymous):

i still don't seem to be able to correctly write it as a prove

OpenStudy (ybarrap):

Yes, so let g(t) be 1 and f(t) stay as f(t). Then the convolution becomes the integral you are trying to take the Laplace of. But now we can use the convolution theorem that says, the Laplace of a convolution is just the product of the Laplace of each individual function. What is the Laplace of "1"? What is the Laplace of f(t) ?

OpenStudy (anonymous):

t/s

OpenStudy (anonymous):

multiplied together = t/s, in this case 1/s

OpenStudy (ybarrap):

If you look at the equation above, with g(t) = 1, the convolution of g(t)*f(t)=1*f(t) = $$\int^t_0 f(u)du$$, which is the integral you want to find the Laplace of? Right?

OpenStudy (anonymous):

yes

OpenStudy (ybarrap):

So finding the Laplace of $$\int^t_0 f(u)du$$ is the same as finding the Laplace of g(t)*f(t). This is the case because both are equal. Do you see that?

OpenStudy (anonymous):

i think so...im also fairly new to the lace transform, so sorry for how slow i am

OpenStudy (ybarrap):

That's ok. It's just important that you see that \(\large \tt g(t)*f(t)= \int^t_0 g(t-u)f(u) du =\int^t_0 f(u)du\), when \(\tt g(t) = 1.\)

OpenStudy (ybarrap):

This is true because g(t-u) = 1, always in the integrand, when g(t)=1.

OpenStudy (ybarrap):

Here is the basic result: $$\frak{L}\tt[f(t)*g(t)]\\ =\frak{L} \tt \left[ \int^t_0 f(u)du\ \right] \tt\\ =F(s)G(s) \\ \text{but}\\ G(s)=\int^\infty_0e^{-st}g(t)dt =\int^\infty_0e^{-st}dt =\frac{1}{s}\\ So~, \frak{L} \tt \left[ \int^t_0 f(u)du\ \right] \tt=\frac{F(s)}{s}\\ $$ Does this make sense?